The diameters of ball bearings are distributed normally. The mean diameter is 96 millimeters and the standard deviation is 6 millimeters. Find the probability that the diameter of a selected bearing is greater than 105 millimeters. Round your answer to four decimal places.
Answer #1

Solution :

Given that,

mean = = 96

standard deviation = = 6

P(x > 105) = 1 - P(x<105 )

= 1 - P[(x -) / < (105-96) /6 ]

= 1 - P(z <1.8 )

Using z table

= 1 -  0.9641

probability= 0.0359